OpenStudy (anonymous):

Calculus help! Find the extreme values of the function and where they occur. f(x)=x-4sqrt(x)

OpenStudy (anonymous):

Find the derivative and set it equal to zero.

OpenStudy (anonymous):

The maximum and minimum are where y' = 0 By the product rule, if y = f(x) g(x), then y' = f(x) g'(x) + f'(x) g(x). So, y' = x^2 e^x + 2x e^x = 0; divide through by e^x, x^2 + 2x = 0; factoring, x (2 + x) = 0, so x = 0 and x = -2. Substitute into the original equation to get the y's, and we obtain (0,0) and (-2, 0.54) as the extremes. Since the function always has a positive slope for x>0, then (0,0) must be a function minimum. Therefore (-2, 0.54) is the function maximum.

OpenStudy (anonymous):

so the derivative would be =1-2x^-1/2?

OpenStudy (experimentx):

occurs at 1-2/sqrt(x) = 0 =>solve for x,

OpenStudy (anonymous):

\[f(x)=x-4\sqrt{x}\] is the equation

OpenStudy (anonymous):

Looks like you are going to have a cusp because you are going to have an imaginary number as one of the values... No negatives in square roots.

OpenStudy (experimentx):

http://www4c.wolframalpha.com/Calculate/MSP/MSP3891a12dab5i5i5822400004i88c35ceba41di2?MSPStoreType=image/gif&s=12&w=320&h=130&cdf=RangeControl at the point where you see following shape